# IDEA: An Inverse Domain Expert Adaptation Based Active DNN IP Protection Method

### v1

@@ VIII Conclusion
We have presented a novel active DNN IP protection method called IDEA. It steganographically encodes distributed model instances with user-specific keys. The authorized user of a protected model instance can unlock its inference performance by submitting test images embedded with a valid key assigned by the model owner. In the event of IP infringement, the owner can verify the authenticity of deployed models and trace the culprit by querying the suspected model with a small set of test samples or decoding the user submitted queries. Extensive experiments conducted across five image classification datasets and four DNN models validated the excellent inference performance on authorized inputs, and corroborated that IDEA-protected models cannot be unlocked by images encoded with incorrect keys even with only a few flipped bits or keys stolen from other legitimate users. This strong uniqueness property is important to assure that the tracking of any redistributed purchased models to its culprit is non-repudiable. The culprit can be traced by decoding the intercepted test images with almost 100% success rate. IDEA is also robust against model transformation and reverse-engineering attacks, and more stealthy, effective, and scalable with networks and dataset sizes compared with existing active protection methods.

### v2

@@ IX Conclusion
We have presented a novel active DNN IP protection method called IDEA. It steganographically encodes distributed model instances with user-specific keys. The authorized user of a protected model instance can unlock its inference performance by submitting test images embedded with a valid key assigned by the model owner. The licensee can be verified through SteganoGAN decoding with almost 100% accuracy, without exposing authorized users’ secret keys. In the event of IP infringement, the owner can verify the authenticity of deployed models and trace the culprit by querying the suspected model with a small set of test samples or decoding the user submitted queries. Extensive experiments conducted across four image classification datasets and four DNN models validated the excellent inference performance on authorized inputs, and corroborated that IDEA-protected models cannot be unlocked by images encoded with incorrect keys even with only a few flipped bits or keys stolen from other legitimate users. This strong uniqueness property is important to assure that the tracking of any redistributed models to its culprit is non-repudiable. IDEA is also robust against model transformation and reverse engineering attacks, and more stealthy, effective, and scalable with networks and dataset sizes compared with existing active protection methods.